Michelle has 4 large storage chests in her basement. Each storage chest has a base of 8 square feet. Three of the storage chests have a height of 2 feet and one storage chest has a height of 3 feet. What is the total volume, in cubic feet, of the four storage chests?
A. 48
B. 72 (the correct answer)
C. 88
D. 96

I need to know how to determine the correct answer of 72.

Answer by mananth(16946)

4 chests base area of each chest = 8 ft^2
3 chests height = 2 feet
Volume of chest = area of base * height
Volume of one chest = 8*2 =16 ft^3
Volume of 3 chests = 16*3 =48 ft^3 ------------------------1
Volume of 4th. Chest = 8 * 3 =24 ft^3-----------------------2
Total volume = 48+24 = 72 ft^3